1.When copper (II) chloride react with sodium nitrate, copper (II) nitrate and sodium chloride are formed. Write the balanced equation for the reaction given above:
B) CuCI2+NaNO3=Cu(NO3)2+NaCI
C) If 15 grams of copper (II) chloride react with 20 grams of sodium nitrate, how much sodium chloride can be formed?
D) What is the limiting react ant and the excess reactant?
E) How many grams of the excess reagent is left over in this reaction?
F)What will be the mass of Cu(NO3)2 produced if the percentage yield is 92%?
1
Expert's answer
2020-11-09T14:09:24-0500

B) CuCI2+2NaNO3=Cu(NO3)2+2NaCI;

C) m(NaCl)=m(CuCI2)*2Mr(NaCl)/Mr(CuCI2)=15*2*94/135=20.89 g

D) CuCI2 - limiting reactant;

NaNO3 - excess reactant

E)mleft(NaNO3)=20 - (15*170/135)=1.1 g

F) Cu(NO3)2=0.92*15*188/135=19.22 g